OpenStudy (narissa):

which number is prime 15,27,43,63,99?

OpenStudy (abdullahm):

What is 15 divisible? 27? 43? 63? 99?

rebeccaxhawaii (rebeccaxhawaii):

prime means the only factors they have is 1 and itself

rebeccaxhawaii (rebeccaxhawaii):

15 isnt prime because 3*5=15

rebeccaxhawaii (rebeccaxhawaii):

27 isnt prime because 3*9=27

rebeccaxhawaii (rebeccaxhawaii):

99 isnt prime because 11*9=99 etc

OpenStudy (abdullahm):

Hint: When you're trying to find factors, start dividing the numbers by 2 as much as you can. Then move to 3, then 5, 7, 11, 13, etc. Teachers usually won't give you huge prime numbers.

OpenStudy (narissa):

9*7=63 so 43 must be prime

OpenStudy (narissa):

Thank you! @AbdullahM @rebeccaxhawaii

OpenStudy (abdullahm):

It was my pleasure!

OpenStudy (abdullahm):

Also, a tip: Don't get confused with squares and primes. 9 is not a prime number because it can be broken down to 3 * 3.